Classify the following triangles are acute obtuse or right 24 11 12 10 30 10
Answer #1

(10) Güven triangle ) 24 А 30 10 B 2 hese (24)? + 102 = 576 +100 = 676 + (30) hence there is not right angle in DABC Hence tA (12) 6 5 B 7 С Hese 2 6)2 +16)2 = 25+36 = 61 #(7)? 15)? +(6)?(7)2 here is not satisfied the pythagos as the sem. le, there(10) acute obtuse

(11) acute obtuse

(12) acute obtuse

Answers...